Accounting Rate of Return

WeCare Clinic is planning on investing in some new echocardiogram equipment that will require an initial outlay of $145,000. The system has an expected life of five years and no expected salvage value. The investment is expected to produce the following net cash flows over its life: $88,000, $76,000, $84,000, $83,000, and $95,000.

Required:

1. Calculate the annual net income for each of the five years.

Net Income
Year 1 $
Year 2 $
Year 3 $
Year 4 $
Year 5 $

2. Calculate the accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16").
%

3. What if a second competing revenue-producing investment has the same initial outlay and salvage value but the following cash flows (in chronological sequence): $95,000, $95,000, $95,000, $88,000, and $26,500? Calculate its accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16").
%
Answer #1

Solution 1:

Annual depreciation = $145,000 / 5 = $29,000

Computation of annual net income
Cash Inflows Depreciation Net Income
Year 1 $88,000.00 $29,000.00 $59,000.00
Year 2 $76,000.00 $29,000.00 $47,000.00
Year 3 $84,000.00 $29,000.00 $55,000.00
Year 4 $83,000.00 $29,000.00 $54,000.00
Year 5 $95,000.00 $29,000.00 $66,000.00

Solution 2:

Average annual income = ($59,000 + $47,000 + $55,000 + $54,000 + $66,000) / 5 = $56,200

Accounting rate of return = Average annual income / Initial investment = $56,200 / $145,000 = 39%

Solution 3:

Computation of annual net income
Cash Inflows Depreciation Net Income
Year 1 $95,000.00 $29,000.00 $66,000.00
Year 2 $95,000.00 $29,000.00 $66,000.00
Year 3 $95,000.00 $29,000.00 $66,000.00
Year 4 $88,000.00 $29,000.00 $59,000.00
Year 5 $26,500.00 $29,000.00 -$2,500.00

Average annual income = ($66,000 + $66,000 + $66,000 + $59,000 - $2,500) / 5 = $50,900

Accounting rate of return = Average annual income / Initial investment = $50,900 / $145,000 = 35%